# Adaptive Multimodal Person Recognition: A Robust Framework for Handling Missing Modalities

### v1

@@ 4.1 CANDOR Dataset
The CANDOR dataset comprises recordings of 1,450 speakers engaged in discussions on various topics, such as politics, family, and COVID-19, while interacting with different individuals they are meeting for the first time. Audio and video are recorded throughout these conversations, capturing the speakers’ speech, facial expressions, and gestures. As depicted in the Figure 2, the dataset includes multiple sessions, with approximately half of the speakers appearing in only one session, while some have participated in up to eight sessions. Additionally, 21 speakers either turned off their cameras, had obstructions in front of them, faced poor lighting conditions, or wore masks, leading to their exclusion from the evaluation.
@@
Ultimately, the dataset includes 1,429 speakers, each with 16 to 50 training utterances, and 8 to 40 validation and test utterances, stored in MP4 format to extract audio, facial, and gesture features. The number of utterances in each set is reported in the Table 1.
4.2 VoxCeleb1 Dataset
We employ the VoxCeleb1 dataset [nagrani2017voxceleb] as a benchmark to develop and evaluate our person identification system. VoxCeleb1 is a large-scale, audio-visual dataset containing over 150,000 utterances from 1,251 celebrities, collected from interview videos on YouTube. In our work, we utilize the official train, validation, and test splits provided with the dataset, which allows for a fair and standardized evaluation. The number of utterances used in each split is detailed in Table 1.
@@ 4.2 VoxCeleb1 Dataset
4.3 Data Augmentation
To enhance the robustness of our speaker recognition model, we employ a multimodal data augmentation strategy that introduces controlled variability into the input features. Each modality, which consists of face, gesture, and voice, is augmented independently using a combination of Gaussian noise, dropout regularization, and feature masking.
Dropout is applied at the feature level, randomly deactivating 20% of neurons, which prevents over-reliance on any specific modality while encouraging feature redundancy. Additionally, feature masking is used to simulate partial or complete modality loss by setting a fraction of the feature dimensions to zero, with a probability of 0.2 per batch.
This technique is particularly useful in handling scenarios where one or more modalities are degraded, missing, or occluded, as it forces the model to rely on the remaining modalities for person identification.
5 Experimental Result
In this study, we developed systems under both Bimodal and Trimodal configurations. For the Bimodal setup, we utilized voice and face modalities sourced from the VoxCeleb1 and CANDOR datasets. In contrast, the Trimodal system was trained solely on the CANDOR dataset and incorporates gesture data in addition to face and voice modalities. Finally, to evaluate the contribution of each component, we conducted an ablation study to analyze the impact of individual modules on overall performance.
Table 2 presents the results of the three encoders consisting of ResNet293, IR101-Adaface, and TimesFormer, evaluated for each modality to assess their Unimodal performance. Each model is fine-tuned using a classifier head on the VoxCeleb1 and CANDOR datasets separately. The results indicate that on the CANDOR dataset the baseline models for voice and face achieved Top-1 accuracies of 97.26% and 97.33%, respectively, significantly outperforming the gesture-based model, which obtained a Top-1 accuracy of 75.4%. This suggests that, in this work, the gesture embedding vector presents greater challenges in person identification compared to voice and face features. However, for the VoxCeleb1 dataset the models for voice and face reached 99.51% and 99.74%, respectively. All the models in this paper were trained in 5 to 8 epochs.
Table 2: Top-1 Accuracy (%) of baseline Unimodal person-identification models on the CANDOR and VoxCeleb1 benchmarks.
Modality CANDOR VoxCeleb1
Face 97.33 99.74
@@ 5 Experimental Result
Voice 97.26 99.51
Utilizing the gesture for identity recognition is relatively uncommon, and this feature tends to be highly session-dependent. Based on the results presented in Table 3, we evaluate the performance of all unimodal models under two conditions. The first scenario occurs when the training and testing data for each speaker are from the same session, and the second scenario occurs when they are from different sessions.
Table 3: Top-1 accuracy (%) in person identification for face, gesture, and voice for speakers with the same train and test sessions and speakers with the different sessions
@@
In the multimodal scenarios, to assess the performance in modality loss conditions, we evaluate the model’s performance in single-modality and Bimodal settings by selectively removing the influence of specific modalities. For single-modality evaluation (e.g., voice-only or face-only), the features corresponding to the other two modalities are set to zero before being passed into the network. This ensures that the model processes only the selected modality while maintaining its multimodal inference structure.
Notably, even when two modalities are removed, the model still runs through the confidence-weighted fusion and mistake correction blocks, ensuring that the final prediction is obtained in a manner consistent with the full multimodal setup. This approach provides a fair and realistic measure of how much each modality independently contributes to person identification. This is particularly useful for testing the model’s adaptability in incomplete modality scenarios.
Table 4: Top-1 and Top-5 Accuracy (%) for three different models based on different training strategies. The first model is trained only on CANDOR and tested on CANDOR (first column). The second model is trained only on VoxCeleb1 and tested on VoxCeleb1 (second column). The third model is trained on a mixture of CANDOR and VoxCeleb1, with the third column showing test results on CANDOR and the fourth column showing test results on VoxCeleb1. In the table, ”C” represents CANDOR, and ”V” refers to VoxCeleb1. The bolded numbers are the best results for each dataset.
@@ 5.1 Bimodal System
An analysis of the Top-1 accuracy of Table 4 shows that the Bimodal system outperforms the Unimodal configuration. Specifically, for the CANDOR dataset, the system achieved an accuracy of 99.17%, which is approximately 2% higher than the Unimodal results reported in Table 2. Furthermore, the system demonstrated robust performance under modality loss conditions due to the use of dedicated modules designed to handle missing inputs. When the face modality was removed, accuracy reached 96.17%, and when the voice modality was removed, accuracy was 96.53%.
For the VoxCeleb1 dataset, the Bimodal system achieved an impressive 99.90% Top-1 accuracy, surpassing the Unimodal results shown in Table 2. Even in scenarios where one modality was absent, the system maintained high performance, achieving 99.28% in the face-only case and 98.51% in the voice-only case.
The proposed system demonstrates strong potential to scale to a larger number of speakers. In the mixed-dataset configuration, the combined training on both CANDOR and VoxCeleb1 led to improved performance for each dataset. Despite the increased complexity introduced by nearly doubling the number of identity classes, which makes the classification task more complex, the larger training set enabled the system to learn more discriminative representations. As a result, it achieved its highest performance across all configurations, reaching 99.92% accuracy on VoxCeleb1 and 99.18% on CANDOR.
Additionally, Top-5 accuracy results for all configurations are reported and can be further examined for a more comprehensive evaluation.
5.2 Trimodal System
The proposed Trimodal system, illustrated in Figure 1, was evaluated exclusively on the CANDOR dataset. This evaluation included a comprehensive assessment of all modality loss scenarios.
@@ 5.2 Trimodal System
Table 5: Top-1 and Top-5 Accuracy (%) of Trimodal System under Different Modality Loss conditions on CANDOR Dataset
@@
Top-1% 96.60 66.73 96.10 99.11 97.14 94.98 99.18
Top-5% 98.44 75.56 98.07 99.57 98.67 98.06 99.58
Table 6: Top-1 accuracy (%) in speaker identification using different modality combinations, reported separately for speakers with a single session, multiple sessions, and overall.
Condition Face Gesture Voice Face+Voice Face+Gesture Gesture+Voice Trimodal #Speakers
Single Session 99.28 99.81 97.79 99.69 99.66 98.63 99.56 679
Multiple Sessions 95.84 57.40 95.62 98.94 96.43 93.96 99.07 750
Overall 96.60 66.73 96.10 99.11 97.14 94.98 99.18 1429
Table 7: Ablation study. Top-1 accuracy (%) for each modality combination, bypassing different modules at a time from the full model. Modules were removed in the order shown. $\Delta_{\text{Top-1}}$ shows the drop in Top-1 accuracy for the Trimodal result compared to the full model. The bolded numbers are the best results.
@@
Full Model 96.6 66.7 96.1 99.1 97.1 95.0 99.2
– w/o Mistake Correction 96.2 65.8 95.8 98.8 96.6 94.1 98.7 0.5-0.5
– w/o Cross-Attention 94.1 62.0 94.0 97.1 93.8 90.2 96.7 2.0-2.0
– w/o Gated Fusion 93.8 61.5 93.6 96.8 93.2 89.8 96.4 0.3-0.3
– w/o Confidence 84.0 51.8 81.9 89.1 85.4 79.0 89.6 6.8-6.8
– w/o Augmentation 82.8 48.5 80.7 87.9 84.1 75.7 86.2 3.4-3.4
Table 8: Top-1 identification accuracy (%) on VoxCeleb1 benchmark. Rows are ordered from lowest to highest accuracy
@@
CNN [nagrani2018seeing] Bimodal 81.0
SSL [niizumi2024masked] Voice 81.2
Adversarial ResNet [chung2019delving] Voice 89.0
VGGnet [yadav2018learning] Voice 89.5
Wav2Vec2 + CNN [anidjar2024harnessing] Voice 90.5
SAIC [cheng2024saic] Voice 96.1
Two-branch Fusion [shah2023speaker] Bimodal 97.2
@@
Table 5 presents a performance comparison of the proposed Trimodal person identification systems. Each row reports the Top-1 and Top-5 accuracies across Unimodal, Bimodal, and Trimodal configurations.
Based on the results, the model achieved a Top-1 accuracy of 99.18% in the Trimodal setup, showing a slight improvement over the 99.11% achieved by the face-voice Bimodal system. This performance gain is attributed to the inclusion of the gesture modality.
Additionally, comparing the face-only configuration (96.6%) with the face-gesture configuration (97.14%) indicates that the gesture modality can enhance performance in visual-only scenarios. However, using gesture as an auxiliary input alongside voice resulted in a performance drop from 96.1% in the voice-only setup to 94.98% in the voice-gesture configuration. These findings suggest that the gesture modality is most beneficial when combined with the face modality, likely due to their shared visual nature and complementary information.
Table 6 presents top-1 speaker identification accuracy across different modality combinations, separated by whether the training and testing samples come from the same or different recording sessions. The results highlight the impact of session variability on speaker recognition performance.
In the Single Session setting, all modalities perform exceptionally well, with face and gesture achieving over 99% accuracy and voice slightly lower at 97.8%. This indicates that under consistent conditions, unimodal cues are highly effective. However, in the Multiple Sessions condition, accuracy drops and most notably for gestures (down to 57.4%), suggesting that motion features are particularly sensitive to session changes.
5.3 Ablation and Comparison
Table 7 presents a comprehensive ablation study that systematically evaluates the contribution of each component to our multimodal person recognition system by progressively removing modules from the whole model. The results demonstrate that all components contribute meaningfully to the system’s performance, with some having more critical roles than others. The mistake correction module shows the smallest impact when removed, causing only a 0.5% drop in Trimodal accuracy, suggesting it provides fine-tuning improvements rather than fundamental functionality. Cross-attention and gated fusion mechanisms prove more important, and contributing approximately 2.8% to the overall performance, indicating their crucial role in effectively combining multimodal information.
However, the most significant performance degradations occur when removing the confidence estimation module (9.6% drop) and data augmentation strategies (13.0% drop), highlighting these as the most critical components of the architecture. Notably, the face modality consistently achieves the highest individual performance across all settings.
@@ 5.3 Ablation and Comparison
Table 8 benchmarks our Bimodal person identification system against previous methods on the standard VoxCeleb1 dataset. based on Table 4 proposed model achieves a Top-1 accuracy of 99.92%, significantly outperforming the previous best result of 97.2%. Additionally, the Unimodal performance of our system reaches 99.39%, further demonstrating its effectiveness even without multi-modal integration.
6 Conclusion
@@
In this paper, we introduced a robust multimodal framework for person identification, integrating face, gesture, and voice modalities through an adaptive and confidence-driven fusion strategy. Our Trimodal and Bimodal systems leverage a unified neural architecture incorporating modality-specific processing, cross-attention mechanisms, gated fusion, and mistake correction modules. This design effectively addresses the challenge of modality loss, significantly outperforming conventional unimodal and late-fusion approaches.
@@
Comprehensive evaluations on the CANDOR dataset, for which we established the first benchmark in this paper, and the established VoxCeleb1 benchmark demonstrate the efficacy of our proposed method. The Trimodal system achieves a top-1 accuracy of 99.18% on CANDOR, while our bimodal configuration attains a remarkable 99.92% accuracy on VoxCeleb1, establishing a new state-of-the-art in multimodal person identification. Ablation studies further confirm the importance of the confidence-weighted fusion and data augmentation strategies in achieving robust performance.

### v2

@@ IV-A CANDOR Dataset
The CANDOR dataset comprises recordings of 1,450 speakers engaged in discussions on various topics, such as politics, family, and COVID-19, while interacting with different individuals they are meeting for the first time. Audio and video are recorded throughout these conversations, capturing the speakers’ speech, facial expressions, and gestures. As illustrated in Figure 2, speaker participation across sessions varies significantly; while approximately half of the participants appear in only a single session, others are featured in as many as eight. We excluded 21 speakers from our evaluation due to unusable video data caused by disabled cameras, physical obstructions, or the wearing of masks.
@@
Ultimately, the dataset includes 1,429 speakers, each with 16 to 50 training utterances, and 8 to 40 validation and test utterances, stored in MP4 format to extract audio, facial, and gesture features. The number of utterances in each set is reported in the Table I.
IV-B VoxCeleb1 Dataset
We employ the VoxCeleb1 dataset [24] as a benchmark to develop and evaluate our person identification system. VoxCeleb1 is a large-scale, audio-visual dataset containing over 150,000 utterances from 1,251 celebrities, collected from interview videos on YouTube. In our work, we utilize the official train, validation, and test splits provided with the dataset, which allows for a fair and standardized evaluation. The number of utterances used in each split is detailed in Table I.
@@ IV-B VoxCeleb1 Dataset
IV-C Data Augmentation
To enhance the robustness of our speaker recognition model, we employ a multimodal data augmentation strategy. Each modality, which consists of face, gesture, and voice, is augmented independently using a combination of Gaussian noise, dropout regularization, and feature masking.
Dropout is applied at the feature level, randomly deactivating 20% of neurons, which prevents over-reliance on any specific modality while encouraging feature redundancy. Additionally, feature masking is used to simulate partial or complete modality loss by setting a fraction of the feature dimensions to zero, with a probability of 0.2 per batch. This technique is particularly useful in handling scenarios where one or more modalities are degraded, missing, or occluded, as it forces the model to rely on the remaining modalities.
To improve robustness against multi-session variations, we integrate Distribution Uncertainty (DSU) augmentation into our model with probability p=0.5 and scaling factor 1.0. DSU models feature distribution uncertainty by estimating batch-level statistics and sampling augmented features during training, effectively simulating domain shifts across recording sessions to enhance generalization [16].
V Experimental Result
In this study, we developed systems under bimodal and trimodal configurations and conducted evaluations across the full range of unimodal, bimodal, and trimodal scenarios. For the bimodal setup, we utilized voice and face modalities sourced from the VoxCeleb1 and CANDOR datasets. In contrast, the trimodal system was trained solely on the CANDOR dataset and incorporates gesture data in addition to face and voice modalities. Finally, to evaluate the contribution of each in the proposed system, we conducted an ablation study. All the models in this paper were trained in 5 to 8 epochs. We employed a cosine learning rate schedule with warmup and AdamW [19] as the optimizer.
Table II presents the results of the three encoders consisting of ResNet293, IR101-Adaface, and TimesFormer, evaluated for each modality to assess their unimodal performance. Each model is fine-tuned using a classifier head on the VoxCeleb1 and CANDOR datasets separately. The results indicate that on the VoxCeleb1 dataset the models for voice and face reached 99.51% and 99.74%, respectively. However, for the CANDOR dataset the baseline models for voice and face achieved Top-1 accuracies of 97.26% and 97.33%, respectively, significantly outperforming the gesture-based model, which obtained a Top-1 accuracy of 75.4%. This suggests that the gesture embedding vector is more challenging to use for person identification than voice and face features, meaning we cannot rely on it as a standalone modality.
TABLE II: Top-1 Accuracy (%) of baseline Unimodal person-identification models on the CANDOR and VoxCeleb1 benchmarks.
Modality VoxCeleb1 CANDOR
Face 99.74 97.33
@@ V Experimental Result
Voice 99.51 97.26
Utilizing gesture for identity recognition is relatively uncommon. This modality tends to be highly session-dependent, as model attempts to capture temporal patterns that are sensitive to camera angles and environmental settings. To reveal this dependency, we evaluate the performance of all unimodal models under two conditions: the first where training and testing data for each speaker are drawn from the same session, and the second where they are drawn from different sessions.
TABLE III: Top-1 accuracy (%) in person identification for face, gesture, and voice for speakers with the same train and test sessions and speakers with the different sessions
@@
The detailed results of the session-based evaluation are presented in Table III. For the gesture modality, the model achieves a Top-1 accuracy of 99.93% in same-session scenarios; however, this performance drops significantly to 68.15% when evaluated across different sessions, whereas face and voice modalities maintain relatively stable accuracy. This suggests that session-specific factors, including camera angle, speaking style, and background setup, significantly influence gesture-based identity features. Furthermore, cognitive and contextual variables, such as the conversation topic, the speaker’s mood, and interpersonal dynamics with a partner, potentially affect the results. Consequently, while gesture features are highly reliable within a single session, their generalization across sessions is limited.
TABLE IV: Top-1 and Top-5 Accuracy (%) for three different models based on different training strategies. The first model is trained only on CANDOR and tested on CANDOR (first column). The second model is trained only on VoxCeleb1 and tested on VoxCeleb1 (second column). The third model is trained on a mixture of CANDOR and VoxCeleb1, with the third column showing test results on CANDOR and the fourth column showing test results on VoxCeleb1. In the table, ”C” represents CANDOR, and ”V” refers to VoxCeleb1. The bolded numbers are the best results for each dataset.
@@ V-A Bimodal System
In the multimodal scenarios, to assess the performance in modality loss conditions, we evaluate the model’s performance in single-modality and bimodal settings by selectively removing the influence of specific modalities. For single-modality evaluation (e.g., voice-only or face-only), the features corresponding to the other modality are set to zero before being passed into the network. This ensures that the model processes only the selected modality while maintaining its multimodal inference structure.
@@
The analysis of the Top-1 accuracy of Table IV shows that the bimodal system outperforms the unimodal configuration. Specifically, for the CANDOR dataset, the system achieved an accuracy of 99.17%, which is approximately 2% higher than the unimodal results reported in Table II. The system also exhibited robustness to modality loss. Accuracy reached $96.17\%$ and $96.53\%$ when facial and voice modalities were respectively removed. These results are slightly lower than unimodal baselines due to the joint loss function, which prioritizes optimizing the overall bimodal framework across multiple tasks.
For the VoxCeleb1 dataset, the bimodal system achieved 99.90% Top-1 accuracy, surpassing the unimodal results shown in Table II. Even in scenarios where one modality was absent, the system achieved 99.28% in the face-only case and 98.51% in the voice-only case.
The proposed system demonstrates strong potential to scale to a larger number of speakers. In the mixed-dataset configuration, the combined training on both CANDOR and VoxCeleb1 led to improved performance for each dataset in bimodal mode. Despite the increased complexity introduced by nearly doubling the number of identity classes, the larger training set enabled the system to learn more discriminative representations. As a result, it achieved its highest performance across all bimodal configurations, reaching 99.92% accuracy on VoxCeleb1 and 99.18% on CANDOR.
V-B Trimodal System
The proposed trimodal system, illustrated in Figure 1, was evaluated exclusively on the CANDOR dataset. This evaluation included a comprehensive assessment of all modality loss scenarios. Notably, even when two modalities are removed, the model still runs through the confidence-weighted fusion and mistake correction blocks, ensuring that the final prediction is obtained in a manner consistent with the full multimodal setup. This approach provides a fair and realistic measure of how much each modality independently contributes to person identification. This is particularly useful for testing the model’s adaptability in incomplete modality scenarios.
@@ V-B Trimodal System
TABLE V: Top-1 and Top-5 Accuracy (%) of Trimodal System under Different Modality Loss conditions on CANDOR Dataset
@@
Top-1% 96.42 73.53 96.13 99.11 97.36 94.97 99.51
Top-5% 98.29 82.23 98.82 99.16 98.58 97.73 99.67
TABLE VI: Top-1 accuracy (%) in speaker identification using different modality combinations, reported separately for speakers with a single session, multiple sessions, and overall.
Condition Face Gesture Voice Face+Voice Face+Gesture Gesture+Voice Trimodal #Speakers #Utterances
Single Session 99.17 99.76 97.52 99.82 99.68 98.73 99.87 679 4,490
Multiple Sessions 95.61 65.76 95.72 98.90 96.67 93.86 99.40 750 14,936
Overall 96.42 73.53 96.13 99.11 97.36 94.97 99.51 1,429 19,426
TABLE VII: Ablation study. Top-1 accuracy (%) for each modality combination, bypassing different modules at a time from the full model. Modules were removed in the order shown. $\Delta_{\text{Top-1}}$ shows the drop in Top-1 accuracy for the Trimodal result compared to the full model. The bolded numbers are the best results.
@@
Full Model 96.42 73.53 96.13 99.11 97.36 94.97 99.51
– w/o Mistake Correction 96.1 72.9 95.7 98.8 97.0 94.1 99.0 0.5-0.5
– w/o Cross-Attention 94.4 70.3 94.1 96.9 94.6 92.3 97.0 2.0-2.0
– w/o Gated Fusion 94.2 70.0 93.7 96.7 94.2 92.1 96.7 0.3-0.3
– w/o Confidence 87.7 63.1 86.9 90.4 89.7 85.8 89.9 6.8-6.8
– w/o Augmentation 84.2 58.6 82.7 87.8 86.1 82.7 86.5 3.4-3.4
Table V presents a performance comparison of the proposed trimodal person identification systems. Each row reports the Top-1 and Top-5 accuracies across unimodal, bimodal, and trimodal evaluation.
Based on the results, the model achieved a Top-1 accuracy of 99.51% in the trimodal setup, showing 0.5% improvement over the 99.11% achieved by the face-voice bimodal scenario. This performance gain is attributed to the inclusion of the gesture modality.
Additionally, comparing the face-only configuration (96.42%) with the face-gesture configuration (97.36%) indicates that the gesture modality can enhance performance in visual scenarios. However, using gesture as an auxiliary input alongside voice resulted in a performance drop from 96.13% in the voice-only setup to 94.97% in the voice-gesture configuration. These findings suggest that the gesture modality is most beneficial when combined with the face modality, likely due to their shared visual nature and complementary information.
Table VI presents top-1 speaker identification accuracy across different modality combinations, separated by whether the training and testing samples come from the same or different recording sessions. The results highlight the impact of session variability on speaker recognition performance.
In the Single Session setting, all modalities perform exceptionally well, with face and gesture achieving over 99% accuracy and voice slightly lower at 97.52%. This indicates that under consistent conditions, unimodal cues are highly effective. However, in the multiple sessions condition, accuracy drops and most notably for gestures (down to 65.76%), suggesting that motion features are particularly sensitive to session changes.
V-C Ablation and Comparison
TABLE VIII: Top-1 identification accuracy (%) on VoxCeleb1 benchmark. Rows are ordered from lowest to highest accuracy
@@ V-C Ablation and Comparison
CNN [23] Bimodal 81.0
SSL [25] Voice 81.2
Adversarial ResNet [6] Voice 89.0
VGGnet [35] Voice 89.5
Wav2Vec2 + CNN [1] Voice 90.5
SAIC [5] Voice 96.1
Two-branch Fusion [31] Bimodal 97.2
@@
To evaluate the contribution of each architectural component to the system’s overall performance, we conducted a systematic ablation study as shown in Table VII.
The most substantial performance drops were observed when removing the Confidence module and Data Augmentation. The exclusion of the Confidence module resulted in a significant decrease of $6.8$ percentage points in trimodal accuracy, highlighting its critical role in dynamically weighting features based on their reliability. Furthermore, the removal of Data Augmentation led to an additional $3.4\%$ drop, with unimodal performance, particularly for gestures, falling to its lowest point ($58.6\%$). These findings suggest that while the fusion logic is vital, the model’s ability to generalize and handle noise is heavily dependent on robust training signals and the ability to evaluate the quality of incoming data.
Finally, the architectural fusion components, specifically Cross-Attention, Gated Fusion, and Mistake Correction, provide the fine-grained refinement necessary for high-precision classification. Cross-Attention proved to be the most influential of these, contributing a $2.0\%$ boost to the trimodal result by capturing complex inter-modality dependencies. Although the individual contributions of Gated Fusion and Mistake Correction appear smaller (drops of $0.3\%$ and $0.5\%$, respectively), they ensure the system’s robustness in bimodal scenarios. For instance, the system maintains high accuracy ($94\%$) even when one modality is missing, proving that these modules successfully mitigate the impact of incomplete data through effective error correction and gated feature selection.
@@
Table VIII benchmarks our Bimodal person identification system against previous methods on the standard VoxCeleb1 dataset. based on Table IV proposed model achieves a Top-1 accuracy of 99.92%, significantly outperforming the previous best result of 97.2%. Additionally, the Unimodal performance of our system reaches 99.39%, further demonstrating its effectiveness even without multi-modal integration.
VI Conclusion
@@
In this paper, we introduced a robust multimodal framework for person identification, integrating face, gesture, and voice modalities through an adaptive and confidence-driven fusion strategy. Our Trimodal and Bimodal systems leverage a unified architecture incorporating modality-specific processing, cross-attention mechanisms, gated fusion, and mistake correction modules. This design effectively addresses the challenge of modality loss, significantly outperforming conventional unimodal and late-fusion approaches.
@@
Comprehensive evaluations on the CANDOR dataset, for which we established the first benchmark in this paper, and the VoxCeleb1 benchmark demonstrate the efficacy of our proposed method. The Trimodal system achieves a top-1 accuracy of 99.51% on CANDOR, while our bimodal configuration attains a remarkable 99.92% accuracy on VoxCeleb1, establishing a new state-of-the-art in multimodal person identification. Ablation studies further confirm the importance of the confidence-weighted fusion and data augmentation strategies in achieving robust performance.